OpenStudy (anonymous):

lim n-> Infinity of summation i=0 to n of (1+i3/n)^(1/2) * 3/n

OpenStudy (amistre64):

looks like reimann to integration

OpenStudy (anonymous):

yes my final answer was 2/3*X^(3/2) from 1to 4

OpenStudy (anonymous):

is that right

OpenStudy (amistre64):

b-a = 3, a = 1, so b = 4 is good

OpenStudy (amistre64):

\[f(a+i\frac{b-a}{n})=(1+x)^{.5}\] the change in x represents x right?

OpenStudy (amistre64):

sqrt(x) is prolly more like it

OpenStudy (amistre64):

\[\int_{1}^{4}x^{.5}dx=\frac{4^{1.5}}{1.5}-\frac{1^{1.5}}{1.5}\]

OpenStudy (anonymous):

okay yes that's what i got i just haven't plugged in the limits to x yet

OpenStudy (anonymous):

thanks for your help would you mind helping me with one more

OpenStudy (amistre64):

if i can remember that far back .. sure :)

OpenStudy (anonymous):

\[d/dx \int\limits_{2x}^{x^2}\sqrt[3]{1+x^4}\]

OpenStudy (amistre64):

in general:\[let:~F(x)=\int_{\alpha(x)}^{\beta(x)}f(t)~dt=F[\beta(x)]-F[\alpha(x)]\] \[\frac d{dx}F(x)=\frac d{dx}\int_{\alpha(x)}^{\beta(x)}f(t)~dt\] \[\frac d{dx}F(x)=\frac d{dx}F[\beta(x)]-F[\alpha(x)]\] \[\frac d{dx}F(x)=f[\beta(x)]\beta'-f[\alpha(x)]\alpha'\]

OpenStudy (amistre64):

\[d/dx \int\limits_{2x}^{x^2}\sqrt[3]{1+t^4}~dt=\sqrt[3]{1+(2x)^4}(2x)'-\sqrt[3]{1+(x^2)^4}(x^2)'\]

OpenStudy (anonymous):

thanks a lot I really appreciate it

OpenStudy (amistre64):

youre welcome